OpenStudy (anonymous):

Completely factor the trinomial or indicate if not factorable: 24b^3+52b^2+24b

jhonyy9 (jhonyy9):

24b*3 +52b*2 +24b /4b 6b*2 +13b +6 6b*2+12b+6+b b*2 +2b +1+b/6 (b+1)*2 +b/6

OpenStudy (anonymous):

= 4b(6b^2 + 13b + 6) = 4b(6b^2 + 9b + 4b + 6) = 4b(3b(2b + 3) + 2(2b + 3)) = 4b(3b + 2)(2b + 3)